# Enhancing Photometric Redshift Catalogs Through Color-Space Analysis: Application to KiDS-Bright Galaxies

### v1

@@ Abstract
We present a method to refine photometric redshift galaxy catalogs by comparing their color-space matching with overlapping spectroscopic calibration data. We focus on cases where photometric redshifts (photo-$z$) are estimated empirically. Identifying galaxies that are poorly represented in spectroscopic data is crucial, as their photo-$z$ may be unreliable due to extrapolation beyond the training sample. Our approach uses a self-organizing map (SOM) to project a multi-dimensional parameter space of magnitudes and colors onto a 2-D manifold, allowing us to analyze the resulting patterns as a function of various galaxy properties. Using SOM, we compare the Kilo-Degree Survey bright galaxy sample (KiDS-Bright), limited to $r<20$ mag, with various spectroscopic samples, including the Galaxy And Mass Assembly (GAMA). Our analysis reveals that GAMA under-represents KiDS-Bright at its faintest ($r\gtrsim 19.5$) and highest-redshift ($z\gtrsim 0.4$) ranges, however no strong trends in color or stellar mass. By incorporating additional spectroscopic data from the SDSS, 2dF, and early DESI, we identify SOM cells where photo-$z$ are estimated suboptimally. We derive a set of SOM-based criteria to refine the photometric sample and improve photo-$z$ statistics. For the KiDS-Bright sample, this improvement is modest: exclusion of the least represented 20% of the sample reduces photo-$z$ scatter by less than 10%. We conclude that GAMA, used for KiDS-Bright photo-$z$ training, is sufficiently representative for reliable redshift estimation across most of the color space. Future spectroscopic data from surveys such as DESI should be better suited for exploiting the full improvement potential of our method.
@@ 1 Introduction
Deep wide-angle imaging of the sky gives access to key probes of cosmology. Among these, weak gravitational lensing by the large-scale structure stands out as it allows us to study the distribution of both dark and luminous matter, and the overall properties of the Universe (Bartelmann & Schneider, 2001; Hoekstra & Jain, 2008). Current ‘Stage-III’ imaging surveys such as the Dark Energy Survey (DES, The Dark Energy Survey Collaboration, 2005), Hyper Suprime-Cam Subaru Strategic Program (HSC, Aihara et al., 2018), and Kilo-Degree Survey (KiDS, Kuijken et al., 2019) have already covered thousands of square degrees and cataloged hundreds of millions of galaxies. The already launched Euclid telescope (Amiaux et al., 2012) and soon-to-be-started Legacy Survey of Space and Time (Ivezić et al., 2019) at the Vera Rubin Observatory will mark the advent of the Stage-IV era when most of the extragalactic sky will be covered and billions of galaxies will be observed.
@@
The constraints on $\Omega_{\mathrm{m}}$ and $\sigma_{8}$ from cosmic shear are degenerate, so extra measurements are needed to separate them. This is in particular possible thanks to the so-called multi-probe approach, where other observables are also used. Among these, the 3$\times$2-point approach has been successfully applied (e.g. Heymans et al., 2021; Abbott et al., 2022; Miyatake et al., 2023). It combines cosmic shear, galaxy-galaxy lensing (GGL; shape distortions of background sources by foreground lenses), and clustering of the lenses.
The two main ways of selecting foreground galaxies for multi-probe analyses are: using spectroscopic data that overlap (at least partly) with the imaging survey; or extracting these galaxies from the photometric sample itself. This former method presents several advantages, such as access to high precision spectroscopic redshift estimates (spec-$z$) and the ability to measure redshift-space distortions and baryon acoustic oscillations from 3-dimensional clustering, strengthening cosmological constraints and breaking degeneracies between cosmological parameters. However, up till now, the overlap of imaging datasets with appropriate wide-angle spectroscopic surveys for cosmic shear analysis and the number densities of the latter have been far from ideal. This leads, for instance, to the GGL signal being sub-dominant with respect to cosmic shear and clustering (e.g. Heymans et al., 2021) and effectively reducing the analysis to 2$\times$2-point if the spectroscopic foreground is used.
In this context, selecting the foreground galaxies directly from the imaging survey is often beneficial, as it offers full-survey coverage, dense sampling, and control over the sample selection function. This comes at the cost, however, of redshift estimate precision: photometric redshift estimates (photo-$z$), derived directly from multi-band imaging datasets, are typically of considerably lower quality than spectroscopic estimates. This in turn allows us to measure only two-dimensional (projected) clustering in redshift bins. In addition, foreground galaxy selection from photometric data risks propagating observational systematic selection effects from the imaging into the lens sample, which can affect the clustering signal and redshift estimates. What is more, such systematics are generally more difficult to compensate for in photometric than in spectroscopic redshift surveys, due to lack of exact information on the radial dimension. When using photometric galaxies for clustering measurements or as lenses in GGL, the quality requirements for their photo-$z$ are more stringent than those for source galaxies. Source galaxies are typically grouped in relatively broad redshift bins, making it crucial to constrain the population redshift distributions accurately. Conversely, for lenses, it is important to have precise knowledge of individual photo-$z$ and their uncertainties.
@@
In this work, we present a method to further calibrate the overlap between spectroscopic training and photometric target data, and apply it to KiDS. In particular, we focus on the ‘KiDS-Bright’ galaxy sample that constitutes one of the foreground photometric datasets used in KiDS for GGL (Brouwer et al., 2021; Georgiou et al., 2021; Burger et al., 2023; Dvornik et al., 2023), the other being luminous red galaxies (Vakili et al., 2019, 2023). Our approach can be used to assess the representativeness of spec-$z$ samples with respect to the photometric ones, and also to improve the photo-$z$ precision of the latter, for instance by discarding those galaxies for which photo-$z$ performance is poor. We choose to perform such a cleaning of galaxies in color-space, using a Self-Organizing Map (SOM, Kohonen, 1982, 2001) by identifying galaxies in the photometric catalog that do not have counterparts in the available spec-$z$ sample.
@@
Using KiDS Data Release 3 (DR3, de Jong et al., 2017) and DR4 (Kuijken et al., 2019), respectively, B18 and B21 selected flux-limited galaxy samples at $r<20$ (mean $z\sim 0.23$) with negligible overall photo-$z$ bias, and scatter $\sigma_{\delta z}\sim 0.018(1+z)$. The resulting KiDS-Bright dataset [^1] Available for download at https://kids.strw.leidenuniv.nl/DR4/brightsample.php. from the latter work covers the full DR4 footprint and contains roughly 1 million galaxies of surface density $\sim 1000$ deg-2. Both the selection of KiDS-Bright galaxies and the derived photo-$z$ were calibrated on the Galaxy And Mass Assembly spectroscopic survey (GAMA, Driver et al., 2011). In its equatorial fields, which are fully covered by KiDS, GAMA offers close to 100%-complete spectroscopic measurements down to $r\sim 19.8$ in SDSS Petrosian magnitude ($r_{\mathrm{Petro}}$). Thanks to the color-independent flux-limited galaxy selection both in GAMA and in KiDS-Bright, B21 were able to obtain high-quality photo-$z$ in the latter not only for red galaxies but also for the blue ones, generally known to perform more poorly in this context. This was done by employing an artificial neural network approach (ANNz2, Sadeh et al., 2016), taking advantage of the very good match between the GAMA training set and the output KiDS-Bright sample. Furthermore, in a recent follow-up William et al. (2023) improved these photo-$z$ further by employing deep-learning methodologies.
@@
At low redshifts, lensing efficiency increases with the depth of the foreground sample. B21 therefore decided to push to slightly fainter limits than allowed by GAMA completeness, taking the risk of including some galaxies not well matched to GAMA spectroscopy. Going slightly deeper than GAMA completeness was possible as that dataset includes a number of ‘filler’ targets at $r_{\mathrm{Petro}}>19.8$ (Baldry et al., 2010); B21 could, therefore, still train photo-$z$ rather robustly at the very faint end of KiDS-Bright.
@@
Even very accurate and precise photo-$z$ may present variations with galaxy properties, such as e.g. color, magnitude, and type. Indeed, B21 confirmed that red galaxies in KiDS-Bright have considerably better photo-$z$ than the blue ones. Some possible $r$-band magnitude dependence in photo-$z$ quality could also be observed at the faint end. Such dependencies often need to be identified and quantified for subsequent applications (e.g. Burger et al., 2023), while this may not always be possible if the spectroscopic calibration data do not fully represent the photometric sample in magnitude or color.
@@
In this work, we develop a method to quantify the completeness of the GAMA spectroscopic calibration data, for use in the photometric selection of KiDS-Bright foreground galaxies. We also show how a color-space comparison can be used to clean up the galaxy sample from objects not meeting a given criterion, for instance, photo-$z$ quality. For that, we use the self-organizing map (SOM): a well-established tool for dimensionality reduction, which allows us to project a multi-dimensional parameter space onto a two-dimensional manifold while maintaining local associations between the objects from the original distribution in higher dimensional space. Unlike supervised machine learning, which uses spectroscopic data as the training dataset here in unsupervised machine learning, we use the photometric sample as the training dataset. Therefore, to avoid confusion, we refer to the spectroscopic data as calibration data.
@@
Our approach borrows from earlier applications in photometric surveys. Masters et al. (2015) utilized a SOM to identify photometric galaxies lacking spectroscopic representation for further observational follow-up, leading to the conception of the Complete Calibration of the Color-Redshift Relation (C3R2) Survey (Masters et al., 2017, 2019), carried out in the context of Euclid (Stanford et al., 2021; Euclid Collaboration et al., 2022), 4MOST (Gruen et al., 2023), and DESI (McCullough et al., 2023). Wright et al. (2020a) developed a SOM-based framework to calibrate redshift distributions for weak lensing shear catalogs, and identify sources that were not represented in spectroscopic calibration data, resulting in a ‘gold’ galaxy selection for weak lensing. This approach has been used extensively for KiDS analyses of cosmic shear (see, e.g., Wright et al., 2020b; Hildebrandt et al., 2021; van den Busch et al., 2022). Furthermore, SOMs have also been utilized in the context of redshift calibration for DES (see, e.g., Buchs et al., 2019; Myles et al., 2021).
@@
In our study, we first show how SOMs can be used to compare color-space coverage between KiDS-Bright and GAMA. We then extend the spectroscopic sample by adding various surveys overlapping with the main KiDS footprint and several external ‘KiDZ’ fields (Wright et al., 2024). As each SOM cell has associated average properties of a given sample (e.g. magnitudes, colors, redshifts), a comparison between, for instance, mean photo-$z$ of KiDS-Bright and mean spec-$z$ of the calibration data can be used to eliminate those photometric galaxies where redshifts are estimated poorly, similarly as done in the KiDS Gold selection (Wright et al., 2020a). Finally, by identifying those SOM cells where spectroscopic representation in KiDS-Bright is still poor, we can select the photometric galaxies for possible future spectroscopic follow-up, similar to the C3R2 survey.
@@
The paper is organized as follows. Section 2 describes the terminologies of the self-organizing maps and their astrophysical application. In Sect. 3, we present the datasets utilized in this work. The projection of SOM and the completeness of GAMA with respect to KiDS is detailed in Sect .4.1. This is followed by SOM comparison of the photometric and spectroscopic data in Sect. 4.2 that allows us to remove the galaxies with worst-constrained photo-$z$s. We conclude and summarize in Sect. 5. In the Appendix, we discuss some further details of galaxy removal based on photo-$z$ performance.
@@ 2 Self organizing maps
A SOM is a form of artificial neural network utilizing unsupervised learning. The main goal of SOM is to transform high-dimensional input into a lower-dimensional representation called a map or grid. This is achieved via the construction of an optimally representative manifold in $n$D, which can be subsequently visualized in two dimensions. The manifold can be planar or toroidal, the latter of which produces a map that wraps horizontally and vertically and is used in this study to avoid edge effects. Additionally, each cell within the map can be rectangular or hexagonal in shape: hexagonal cells being generally preferred, as they allow metrics (such as cell distance) to be computed between more neighbor cells (six compared to four for rectangular cells), thereby providing a finer and more detailed mapping of the input space. Thus, this study uses a toroidal map with hexagonal cells to enhance data representation and analysis.
Figure 1: Illustration of the projection of the $X$-input vector on a two-dimensional SOM, based from figure 1 Fortela et al. (2020). The red dot shows the best-matching unit.
The SOM consists of nodes (or neurons) arranged in a two-dimensional lattice, as shown in Fig. 1. Each node in the map is associated with a weight vector to represent the input patterns. At each training epoch, the algorithm computes the Euclidean distances between the input vector and the weight vectors of all neurons on the map. The node with the weight vector closest to the input vector (i.e. with the smallest Euclidean distance) is identified as the best-matching unit (BMU). The weight vector of this cell (and a subset of the map within the immediate vicinity of the BMU), are adjusted accordingly. During the training process, the learning rate controls the step size of weight adjustments, and the cell shape defines the arrangement and connectivity of the nodes in the map, and the neighborhood function determines the extent of influence a BMU has on its neighboring nodes. The most commonly used neighborhood function is a Gaussian, which assigns weights to each node decreasing with its distance from the BMU according to the bell curve. As the training progresses, the gradual reduction of the radius of the neighborhood function focuses on smaller regions of the map, promoting finer adjustments and improving convergence.
@@
SOMs have been shown to be particularly useful in calibrating redshift distributions and accessing color space match in the photometric and spectroscopic samples (Sánchez et al., 2020; Wright et al., 2020b; Hildebrandt et al., 2021; Razim et al., 2021; Stölzner et al., 2023; Jafariyazani et al., 2024). This is possible because, in addition to photometric data, features from the calibration spectroscopic sample can also be projected onto the SOM. Thanks to this, one can directly identify the cells occupied by both photometric and spectroscopic samples with similar properties.
The cells showing significant deviations between photometric and spectroscopic redshifts indicate the types of galaxies for which the derived photo-$z$ are estimated poorly or the spec-$z$ is not well sampled in the cell. Identifying and removing such cells can help to refine the sample, or be used to single out potential candidates for spectroscopic follow-up to improve the parameter constraints. Therefore, the SOM can be utilized not only to identify a comprehensive representative calibration sample but also to discern clustering among similar galaxies within the feature space, based on their photometric data.
In this paper, we use the SOMOCLU [^2] https://somoclu.readthedocs.io/en/stable/download.html (Wittek et al., 2017) Python package for training SOMs on large data sets. SOMs can be computationally intensive to train, and SOMOCLU is a highly efficient, parallel, and distributed algorithm to train such maps.

### v2

@@ Abstract
Aims. We present a method for refining photometric redshift galaxy catalogs based on a comparison of their color-space matching with overlapping spectroscopic calibration data. We focus on cases where photometric redshifts (photo-$z$) are estimated empirically. Identifying galaxies that are poorly represented in spectroscopic data is crucial, as their photo-$z$ may be unreliable due to extrapolation beyond the training sample.
Methods. Our approach uses a self-organizing map (SOM) to project a multidimensional parameter space of magnitudes and colors onto a 2D manifold, allowing us to analyze the resulting patterns as a function of various galaxy properties. Using SOM, we compared the Kilo-Degree Survey’s bright galaxy sample (KiDS-Bright), limited to $r<20$ mag, with various spectroscopic samples, including the Galaxy And Mass Assembly (GAMA).
Results. Our analysis reveals that GAMA tends to underrepresent KiDS-Bright at its faintest ($r\gtrsim 19.5$) and highest-redshift ($z\gtrsim 0.4$) ranges; however, no strong trends are seen in terms of color or stellar mass. By incorporating additional spectroscopic data from the SDSS, 2dF, and early DESI, we identified SOM cells where the photo-$z$ values are estimated suboptimally. We derived a set of SOM-based criteria to refine the photometric sample and improve photo-$z$ statistics. For the KiDS-Bright sample, this improvement is modest, namely, it excludes the least represented 20% of the sample reduces photo-$z$ scatter by less than 10%.
Conclusions. We conclude that GAMA, used for KiDS-Bright photo-$z$ training, is sufficiently representative for reliable redshift estimation across most of the color space. Future spectroscopic data from surveys such as DESI should be better suited for exploiting the full improvement potential of our method.
Key Words.:
@@ 1 Introduction
Deep wide-angle imaging of the sky gives access to key probes of cosmology. Among these, weak gravitational lensing by the large-scale structure stands out as it allows us to study the distribution of both dark and luminous matter, as well as the overall properties of the Universe (Bartelmann & Schneider, 2001; Hoekstra & Jain, 2008). Current ‘Stage-III’ imaging surveys such as the Dark Energy Survey (DES, The Dark Energy Survey Collaboration, 2005), Hyper Suprime-Cam Subaru Strategic Program (HSC, Aihara et al., 2018), and Kilo-Degree Survey (KiDS, Kuijken et al., 2019) have already covered thousands of square degrees and cataloged hundreds of millions of galaxies. The operational Euclid telescope (Amiaux et al., 2012) and soon-to-be-launched Legacy Survey of Space and Time (Ivezić et al., 2019) at Vera Rubin Observatory will mark the advent of the Stage-IV era, when most of the extragalactic sky will be covered, providing observations of billions of galaxies.
@@
The constraints on $\Omega_{\mathrm{m}}$ and $\sigma_{8}$ from cosmic shear are degenerate, so extra measurements are needed to separate them. This is in particular possible thanks to the so-called multiprobe approach, where other observables are also used. Among these, the 3$\times$2-point approach has been successfully applied (e.g., Heymans et al., 2021; Abbott et al., 2022; Miyatake et al., 2023). It combines cosmic shear, galaxy-galaxy lensing (GGL; shape distortions of background sources by foreground lenses), and clustering of the lenses.
The two main ways of selecting foreground galaxies for multiprobe analyses are either to use spectroscopic data that overlap (at least partly) with the imaging survey or by extracting these galaxies from the photometric sample itself. This former method presents several advantages, such as access to high-precision spectroscopic redshift estimates (spec-$z$) and the ability to measure redshift-space distortions and baryon acoustic oscillations from 3D clustering. This allows us to strengthen the cosmological constraints and break degeneracies between cosmological parameters. However, until now, the overlap of imaging datasets with appropriate wide-angle spectroscopic surveys for cosmic shear analysis and the number densities of the latter have been far from ideal. This leads, for instance, to the GGL signal being sub-dominant with respect to cosmic shear and clustering (e.g., Heymans et al., 2021), effectively reducing the analysis to 2$\times$2-point if the spectroscopic foreground is used.
In this context, selecting the foreground galaxies directly from the imaging survey is often beneficial, as it offers full-survey coverage, dense sampling, and control over the sample selection function. This comes at the cost, however, of redshift estimate precision: photometric redshift estimates (photo-$z$), derived directly from multiband imaging datasets, are typically of considerably lower quality than spectroscopic estimates. This, in turn, allows us to measure only two-dimensional (projected) clustering in redshift bins. In addition, foreground galaxy selection from photometric data risks propagating observational systematic selection effects from the imaging into the lens sample, which can affect the clustering signal and redshift estimates. Furthermore, such systematics are generally more difficult to compensate for in photometric than in spectroscopic redshift surveys, due to a lack of precise information on the radial dimension. When using photometric galaxies for clustering measurements or as lenses in GGL, the quality requirements for their photo-$z$ are more stringent than those for source galaxies. Source galaxies are typically grouped in relatively broad redshift bins, making it crucial to constrain the population redshift distributions accurately. Conversely, for lenses, it is important to have precise knowledge of individual photo-$z$ and their uncertainties.
@@
In this work, we present a method to further calibrate the overlap between spectroscopic training and photometric target data and apply it to KiDS. In particular, we focus on the ‘KiDS-Bright’ galaxy sample, which constitutes one of the foreground photometric datasets used in KiDS for GGL (Brouwer et al., 2021; Georgiou et al., 2021; Burger et al., 2023; Dvornik et al., 2023), the other being luminous red galaxies (Vakili et al., 2019, 2023). Our approach can be used to assess the representativeness of spec-$z$ samples with respect to the photometric ones and also to improve the photo-$z$ precision of the latter; for instance by discarding those galaxies for which photo-$z$ performance is poor. We chose to perform such cleaning of galaxies in color-space, using a self-organizing map (SOM, Kohonen, 1982, 2001) by identifying galaxies in the photometric catalog that do not have counterparts in the available spec-$z$ sample.
@@
Using KiDS Data Release 3 (DR3, de Jong et al., 2017) and DR4 (Kuijken et al., 2019), respectively, B18 and B21 selected flux-limited galaxy samples at $r<20$ (mean $z\sim 0.23$) with a negligible overall photo-$z$ bias and a scatter of $\sigma_{\delta z}\sim 0.018(1+z)$. The resulting KiDS-Bright dataset [^1] Available for download at https://kids.strw.leidenuniv.nl/DR4/brightsample.php. from the latter work covers the full DR4 footprint and contains roughly 1 million galaxies of surface density $\sim 1000$ deg-2. Both the selection of KiDS-Bright galaxies and the derived photo-$z$ were calibrated on the Galaxy And Mass Assembly spectroscopic survey (GAMA, Driver et al., 2011). For its equatorial fields (which are fully covered by KiDS) GAMA offers close to 100% complete spectroscopic measurements down to $r\sim 19.8$ in SDSS Petrosian magnitude ($r_{\mathrm{Petro}}$). Thanks to the color-independent flux-limited galaxy selection both in GAMA and in KiDS-Bright, B21 were able to obtain high-quality photo-$z$ in the latter not only for red galaxies but also for the blue ones, generally known to perform more poorly in this context. This was done by employing an artificial neural network approach (ANNz2, Sadeh et al., 2016), taking advantage of the very good match between the GAMA training set and the output KiDS-Bright sample. Furthermore, in a recent follow-up William et al. (2023) and John William et al (in prep.) improved these photo-$z$ further (reducing scatter by $\sim 20\%$) by employing deep-learning methodologies.
@@
At low redshifts, the lensing efficiency increases with the depth of the foreground sample. Therefore B21 chose to push to slightly fainter limits than allowed by GAMA completeness, taking the risk of including some galaxies not well matched to GAMA spectroscopy. Going slightly deeper than GAMA completeness was possible as that dataset includes a number of ”filler’” targets at $r_{\mathrm{Petro}}>19.8$ (Baldry et al., 2010). Therefore, B21 could still train photo-$z$ rather robustly at the very faint end of KiDS-Bright.
@@
Even very accurate and precise photo-$z$ values may present variations with galaxy properties, such as color, magnitude, and type. Indeed, B21 confirmed that red galaxies in KiDS-Bright have considerably better photo-$z$ than the blue ones. Some possible $r$-band magnitude dependence in photo-$z$ quality could also be observed at the faint end. Such dependencies often need to be identified and quantified for subsequent applications (e.g., Burger et al., 2023), while this may not always be possible if the spectroscopic calibration data do not fully represent the photometric sample in terms of magnitude or color.
@@
In this work, we develop a method to quantify the completeness of the GAMA spectroscopic calibration data, for use in the photometric selection of KiDS-Bright foreground galaxies. We also show how a color-space comparison can be used to clean up the galaxy sample and remove objects that do not meet a given criterion, for instance, the photo-$z$ quality. For that purpose, we used the self-organizing map (SOM): a well-established tool for dimensionality reduction that allows us to project a multidimensional parameter space onto a two-dimensional (2D) manifold, while maintaining local associations between the objects from the original distribution in higher dimensional space. Unlike supervised machine learning, which uses spectroscopic data as the training dataset here in unsupervised machine learning, we used the photometric sample as the training dataset. Therefore, to avoid confusion, we refer to the spectroscopic data as the ”calibration data.”
@@
Our approach follows on earlier applications in photometric surveys. Masters et al. (2015) utilized a SOM to identify photometric galaxies lacking spectroscopic representation for further observational follow-up, leading to the conception of the Complete Calibration of the Color-Redshift Relation (C3R2) Survey (Masters et al., 2017, 2019), carried out in the context of Euclid (Stanford et al., 2021; Euclid Collaboration et al., 2022), 4-meter Multi-Object Spectrograph Telescope (4MOST, Gruen et al., 2023), and Dark Energy Spectroscopic Instrument (DESI, McCullough et al., 2024). Wright et al. (2020a) developed a SOM-based framework to calibrate redshift distributions for weak lensing shear catalogs, and identify sources that were not represented in spectroscopic calibration data, resulting in a ”gold” galaxy selection for weak lensing. This approach has been used extensively for KiDS analyses of cosmic shear (see, e.g., Wright et al., 2020b; Hildebrandt et al., 2021; van den Busch et al., 2022). Furthermore, SOMs have also been utilized in the context of redshift calibration for DES (see, e.g., Buchs et al., 2019; Myles et al., 2021).
@@
In our study, we first show how SOMs can be used to compare color-space coverage between KiDS-Bright and GAMA. We then extended the spectroscopic sample by adding various surveys overlapping with the main KiDS footprint and several external ‘KiDZ’ fields (Wright et al., 2024). As each SOM cell has associated average properties of a given sample (e.g., magnitudes, colors, redshifts), a comparison between, for instance, the mean photo-$z$ of KiDS-Bright and mean spec-$z$ of the calibration data can be used to eliminate those photometric galaxies where redshifts have been estimated poorly, similarly to what was done in the KiDS Gold selection (Wright et al., 2020a). Finally, by identifying those SOM cells where the spectroscopic representation in KiDS-Bright is still poor, we can select the photometric galaxies for possible future spectroscopic follow-ups, similar to the C3R2 survey.
@@
The paper is organized as follows. Section 2 describes the terminologies of the self-organizing maps and their astrophysical application. In Sect. 3, we present the datasets utilized in this work. The projection of SOM and the completeness of GAMA with respect to KiDS is detailed in Sect. 4.1. This is followed by SOM comparison of the photometric and spectroscopic data in Sect. 4.2 that allows us to remove the galaxies with worst-constrained photo-$z$s. We present our conclusions and summary in Sect. 5. In the appendix, we discuss some further details of galaxy removal based on photo-$z$ performance.
@@ 2 Self organizing maps
A SOM is a form of artificial neural network utilizing unsupervised learning. The main goal of SOM is to transform high-dimensional input into a lower dimensional representation called a map or grid. This is achieved via the construction of an optimally representative manifold in $n$D, which can be subsequently visualized in two dimensions. The manifold can be planar or toroidal, the latter of which produces a map that wraps horizontally and vertically and is used in this study to avoid edge effects. Additionally, each cell within the map can be rectangular or hexagonal in shape: hexagonal cells being generally preferred, as they allow metrics (such as cell distance) to be computed between more neighbor cells (six compared to four for rectangular cells), thereby providing a finer and more detailed mapping of the input space. Thus, this study uses a toroidal map with hexagonal cells to enhance data representation and analysis.
Figure 1: Illustration of the projection of the $X$-input vector on a 2D SOM, based on Fig. 1 from Fortela et al. (2020). The red dot shows the best-matching unit.
The SOM consists of nodes (or neurons) arranged in a 2D lattice, as shown in Fig. 1. Each node in the map is associated with a weight vector to represent the input patterns. At each training epoch, the algorithm computes the Euclidean distances between the input vector and the weight vectors of all neurons on the map. The node with the weight vector closest to the input vector (i.e., with the smallest Euclidean distance) is identified as the best-matching unit (BMU). The weight vector of this cell (and a subset of the map within the immediate vicinity of the BMU), are adjusted accordingly. During the training process, the learning rate controls the step size of weight adjustments and the cell shape defines the arrangement and connectivity of the nodes in the map, and the neighborhood function determines the extent of influence a BMU has on its neighboring nodes. The most commonly used neighborhood function is a Gaussian, which assigns weights to each node that are decreasing with its distance from the BMU according to the bell curve. As the training progresses, the gradual reduction of the radius of the neighborhood function focuses on smaller regions of the map, promoting finer adjustments and improving convergence.
@@
SOMs have been shown to be particularly useful in calibrating redshift distributions and accessing color space match in the photometric and spectroscopic samples (Sánchez et al., 2020; Wright et al., 2020b; Hildebrandt et al., 2021; Razim et al., 2021; Stölzner et al., 2023; Jafariyazani et al., 2024). This is possible because, in addition to photometric data, features from the calibration spectroscopic sample can also be projected onto the SOM. Thanks to this, we can directly identify the cells occupied by both photometric and spectroscopic samples with similar properties.
The cells showing significant deviations between photometric and spectroscopic redshifts indicate the types of galaxies for which the derived photo-$z$ are estimated poorly or the spec-$z$ is not well sampled in the cell. Identifying and removing such cells can help to refine the sample. They may also be used to single out potential candidates for spectroscopic follow-up to improve the parameter constraints. Therefore, the SOM can be utilized not only to identify a comprehensive representative calibration sample but also to discern clustering among similar galaxies within the feature space, based on their photometric data.
In this work, we use the SOMOCLU [^2] https://somoclu.readthedocs.io/en/stable/download.html (Wittek et al., 2017) Python package for training SOMs on large data sets. We note that SOMs can be computationally intensive to train and SOMOCLU is a highly efficient, parallel, and distributed algorithm to train such maps.